The sum of two numbers is 20. Twice the first number plus four times the second is 64. Find the numbers

Answer by math_helper(2461)


The two equations for the given conditions are:
n + m = 20 (1)
2n + 4m = 64 (2)
(2) --> n + 2m = 32
Subtract (1) from this to get m:
n + 2m = 32
- n - m = -20
------------------
(you do the rest here)
then plug in that value into (1) or (2) and solve for n.

----
Don't forget to check your answer